The marketing department of Graber Corporation has submitted the following sales forecast for the upcoming fiscal year.

Budgeted units sales: first quarter 17,900, second quarter 16,900, third quarter 15,900, forth quarter 16,900

The selling price of the company’s product is $31.00 per unit. Management expects to collect 85% of sales in the quarter in which the sales are made and 10% in the following quarter, and 5% of sales are expected to be uncollectible. The beginning balance of accounts receivable, all of which is expected to be collected in the first quarter, is $85,000.

The company expects to start the first quarter with 5,100 units in finished goods inventory. Management desires an ending finished goods inventory in each quarter equal to 20% of the next quarter’s budgeted sales. The desired ending finished goods inventory for the fourth quarter is 5,300 units.

Compute the company’s total sales.

Total sales: 1st quarter ?, 2nd quarter ?, 3rd quarter ?, 4th quarter ?, year $?
Answer #1

1st quarter sales = 17900 * 31$ = $554900

2nd quarter sales = 16900 * 31$ =$523900

3rd quarter sales = 15900 * 31$ = $492900

4th quarter sales = 16900 * 31$ =$523900

Total sales = $2095600

Note - I think maybe the question is incomplete.only sales is asked.